# Photonuclear treatment for spent fuel radiotoxicity reduction: a case study investigation on minor actinides

### v1

@@ Abstract
The management of Spent Nuclear Fuel (SNF) is one of the main challenges in the decommissioning of nuclear power plants. Thermal reactors, such as Light Water Reactors (LWRs), produce significant amounts of minor actinides (MAs) such as Americium, Curium, and Neptunium, which are key contributors to the long-term radiotoxicity and decay heat in SNF. Currently, the long term widely accepted solution is the geological disposal. At the same time, advanced technologies like Partitioning and Transmutation (P&T) offer promising solutions to reduce SNF long-term radiotoxicity. While most transmutation strategies rely on neutron fluxes, in this study the adoption of photon beam to induce photonuclear reactions in SNF is investigated, without depending on neutron based systems. In particular, the study focuses on the probability of inducing transmutations and fissions on MAs, by leveraging the Giant Dipole Resonance (GDR) region of photonuclear interactions. In the investigated case study, the effect of a photon driven transmutation of minor actinides present in a spent fuel from SMR technology was evaluated. This approach offers a novel solution to the challenges of nuclear waste management which may become an alternative path in treating the radiotoxicity of minor actinides with the adoption of high enough photon fluxes.
@@
In 2020, the International Atomic Energy Agency (IAEA) reported that 172 Nuclear Power Plants (NPPs) have been permanently shut down and 20 have been fully decommissioned [1]. During the decommissioning and disposal phases, the main concern must be on the management of Spent Nuclear Fuel (SNF) due to its high and long-term radiotoxicity. The composition of SNF varies depending on the reactor design and the fuel burning. Thermal reactors, which operate using a slower neutron spectrum, tend to generate more minor actinides (such as Americium, Curium, and Neptunium) than reactors based on a fast neutron spectrum. Indeed, at thermal neutron energies, the likelihood of neutron absorption is higher, leading to the production of transuranic elements. As a result, the composition of SNF from these reactors typically consists of approximately 94.5% uranium, 4.3% fission products, 1% plutonium (Pu), and 0.2% minor actinides (MAs) [2]111Ratios between elements in SNF depends on the burnup level at the discharge of the reactor. Reported values are representative of SNF at 50 MWd/kg..
@@
Regarding transmutation strategies, the vast majority of designs discussed in the literature focus on neutron-based transmutation, which rely on high neutron fluxes along with a fast neutron spectrum. This creates a scenario where the main driver for the consumption of minor actinides (i.e., neutrons) also contributes to their production. Consequently, there are no viable alternatives for SNF generated by thermal reactors. This limitation has prompted researchers to investigate non-neutron transmutation strategies, including facilities that generate high-energy particles other than neutrons. Recent advancements in photonics and ion accelerators present promising opportunities, enabling the production of energetic non-neutron particles at national lab user facilities or university accelerators.
@@
Bremsstrahlung is the simplest and most widely used method for generating high photon fluxes. This process, induced by the deceleration of intense electron beams in high-Z materials like tungsten, lead, tantalum, or depleted uranium, generates photons with a continuous energy distribution. Materials with higher proton numbers are particularly effective as converters due to their lower critical energy thresholds. Bremsstrahlung facilities are invaluable for generating intense photon beams with straightforward technology, finding applications in experiments such as the search for fission isomers and the production of radioactive ion beams. Despite these advantages, Bremsstrahlung sources have significant limitations, including broad energy spectra, large background signals, and lack of polarization. However, these drawbacks are negligible in programs where the focus is on beam intensity rather than energy precision [19].
In contrast, recent advancements have enabled the production of high-intensity, quasi-monoenergetic gamma beams through Compton backscattering. This process, illustrated in Fig. 1, involves the interaction of laser photons (energy E=h) with GeV-level accelerated electrons: the photons gain energy from the relativistic recoil interaction, with the resulting energy scaling quadratically with the Lorentz factor of the electrons. In this manner, gamma-ray beams in the GDR energy region (10–20 MeV) can be achieved starting with laser photons of relatively low energy. The ability to produce photon beams with a narrow energy bandwidth, strong polarization, and a high degree of tuning makes this method particularly attractive for experiments requiring precise energy control, e.g. for precise evaluation of the GDR region cross section of target nuclei.
@@
Based on the above considerations, in this work a methodology to investigate the photon-driven transmutation of key MAs is developed, having focus on the fundamental physical processes involved in the MAs transmutation and proposing a new strategy for radiotoxicity reduction without relying on neutrons as the primary driver. Specifically, we explore photonuclear interactions for inducing transmutation and even fission by leveraging the Giant Dipole Resonance (GDR) in nuclei. The focus is on photo-absorption reactions, such as photocapture and photofission, triggered by high-energy photons, that have the maximum probability of occurring within the GDR region. To have a complete indication of the system’s behaviour concerning the source, both photons emitted by Bremsstrahlung and LCB are considered. The considered MAs are a subset of the radioactive waste produced by a fuel assembly from the NuScale reactor design [20], one of the most advanced and commercially ready SMR technologies. Specifically, the content of actinides present at the reactor discharge is considered separable into two groups: elements that can be recycled for power generation purposes and those that cannot. The former, such as U and Pu, for example, are not considered to be subject to irradiation, since they can be used in so-called MOX fuels; the latter, such as Np, Am and Cm, which are the main contributors to radiotoxicity, are subject to irradiation. Due to the transmutations that can occur on Np, Am and Cm, this analysis considers the build-up of up to 53 nuclides.
@@ 2.1 Cross section data
Photo-nuclear reactions are fundamental processes in which atomic nuclei interact with high-energy photons, typically high-energy X-rays, resulting in various nuclear transformations. These reactions include photo-disintegration, where a nucleus absorbs a photon and subsequently emits nucleons such as neutrons or protons, and photo-excitation, where the nucleus is excited to a higher energy state [19]. These latter two phenomena appear to have different probability of occurrence depending on the energy of the impinging photon on the nucleus. This probability, encoded in the the photonuclear absorption cross section, can be divided in three main regions[21], depending on the photon energy. The GDR region ( 50 MeV), the Quasideuteron region ( 200 MeV) and the Photomeson region ( 200 MeV). This region is interesting for requiring photon energies lower than 50 MeV and exhibiting high cross sections due to resonant effects. Also, for MAs, this region begins at approximately 5 MeV. More details about the photofission mechanism can be found in Balabanski et al.[19].
@@
- | All 53 nuclides | - | |
- | 232,237,239-240U, 232-236,238-244Np, 232-237,242-244Pu, 234-240,242-245Am, 240-245Cm | 238-241Pu, 233-236,238U, 237Np, 241Am | |
- | 232-233,240-244Np, 232-235Pu, 234-239,245Am | 232-240U, 234-239Np, 236-244Pu, 240-244Am, 240-245Cm | |
- | 232-233,240-244Np, 232-235Pu, 234-239,245Am | 232-240U, 234-239Np, 236-244Pu, 240-244Am, 240-245Cm | |
- | 232-233,240-244Np, 232-235Pu, 234-239,245Am | 232-240U, 234-239Np, 236-244Pu, 240-244Am, 240-245Cm |
Finally, this work relies on a diverse array of nuclear data sources, given the lack of a single comprehensive database containing all relevant reaction information. The considered libraries are JENDL-5 [22], TENDL-2021 [23] and ENDF/B-VIII.0 [24]. Table 1 classifies the cross sections for each considered reactions for the 53 isotopes. It is evident that the data were preferably taken from ENDF/B-VIII.0 and, when necessary, were completed with data taken from the other two libraries.
@@ 2.2 Bateman equation solution
The information of the cross section alone is not sufficient to evaluate the transmutation capability of photons against neutrons. To do so, it is also crucial to compare the absolute values of reactions such as and for neutrons in the energy range above 0.5 MeV (see Figure 2). When neutron and photon fluxes are of equal intensity, the probability of neutron interactions is approximately three times higher than that of photon interactions. However, the actual intensities of these two fluxes differ significantly. The neutron flux is generated as secondary particles during photonuclear interactions and fission, while the photon flux originates from an accelerator, which produces a known and controllable beam. A more complete discussion about fluxes will be faced in Section 3, showing how to extract volume averaged energy fluxes (i.e., ) from the system.
@@ 2.3 Radiotoxicity evaluation
Contents of radioactive isotopes in spent fuel can be assessed in terms of mass, activity or radiotoxicity. The latter figure of merit gives quantitative information on the radiobiological influence on the human body. Hazard from minor actinides is mainly due to alpha decay, which having long half-time, require strategies for their disposal or treatment.
@@
where is the irradiated volume in cm3, are the dose coefficients, tabulated within the ICRP 2012 report [25] and is the outcome concentration of the Bateman equations in at/cm3.
The system radiotoxicity is compared with the free evolution radiotoxicity , namely the Sievert emitted if the same mass of minor actinides used as the initial condition for this calculation were to evolve without any particle flux acting as a driver. This concentration evolution has been evaluated with a new Serpent simulation, adopting the decay step mode, therefore obtaining the concentrations :
Both the quantities and have been normalised with respect to the uranium ore radiotoxicity , calculated as
@@ 3 Case study
The spent fuel inventory is primarily influenced by both the reactor design and the nominal power. The reactor design affects the composition of fresh fuel, including the types of enrichment and fissile materials present, as well as the evolution of the neutron energy spectrum. For instance, a reactor operating with a thermal spectrum tends to burn odd nuclei more quickly, while a reactor with a fast spectrum can effectively burn both odd and even nuclei. The nominal power, on the other hand, determines the rate at which the fuel is consumed, providing insight into the number of minor actinides generated through neutron captures over the reactor lifetime. The figure of merit that encapsulates fuel history is burnup (in MWd/kgU), which is proportional to the total energy produced by the reactor.
@@ 3.1 Initial condition: spent fuel composition
The initial conditions based on the minor actinides present in the fuel can be extracted, as shown in Table 2. It is important to note that uranium and plutonium were deliberately excluded from this calculation, as their reuse in MOX fuels is widely expected in the literature[28]. The actinides considered in this work suffers from being not viable as nuclear fuel at present due to their poor neutronic performances, moreover they are responsible for a significant radiotoxicity level (i.e., 241Am and 237Np).
@@ 3.2 Flux evaluation in different scenarios
The interaction between a particle beam and the spent fuel is not trivial due to the high number of reaction channels available. Since this work uses a lumped approach, each isotope transmutation is driven by an average reaction rate, depending both by the microscopic cross section as well as by the average flux value. To have a precise estimation of the volume-averaged particle flux in a spent fuel sample irradiated by a particle beam, a set of numerical simulations are performed using the FLUKA general purpose MC code for particle transport (version 2024.1.0) [29].
The irradiated sample is a cylinder having radius equal to 1 cm and height equal to 4 cm. This is chosen considering the attenuation of the primary beam and the secondary particles produced inside the spent fuel: the 12 cm3 cylinder delimits the zone where both the primary beam intensity and the secondary particle flux are reduced by a factor 1000. The volume is filled with a fictitious oxide material MO2, where M stands for the isotopic mix shown in Tab. 2. The density has been set equal to 10.2 g/cm3, in the same order of magnitude of common oxides in nuclear field.
@@
The fluxes calculated by FLUKA are known despite a normalization factor F333For example, the FLUKA simulation evaluates the photon flux in the unit of , dependent on the number of primaries per second considered. Since Simulation C adopts a different primary than Simulations A and B, for this work the normalization factor was chosen by constraining the number of within the spent fuel sample.
Namely, for each simulation:
Where is the volume-averaged flux. The value of Fs can also be adopted to normalize the neutron flux . This work considers four different flux scenarios values, i.e. 1E+17, 1E+18, 1E+19, 1E+20 . These values of flux are taken for reference from a similar work made by Matsumoto et al.[17].
Figure 6 reports the scored photon and neutron fluxes in the spent fuel sample in the different simulations, after the normalization of the photon yield to 1E+18 . It is interesting to note few differences in fluxes arising from Bremsstrahlung for diverse Tantalum thicknesses (cases A) and for no Tantalum converter (case B), whereas an LCB source (case C) is able to reproduce quasi-monoenergetic photons in the target volume and a more intense neutron field within the spent fuel.

### v2

@@ Abstract
The management of Spent Nuclear Fuel (SNF) is one of the main challenges in the decommissioning of nuclear power plants. Thermal reactors, such as Light Water Reactors (LWRs), produce significant amounts of minor actinides (MAs) such as Americium, Curium, and Neptunium, which are key contributors to the long-term radiotoxicity and decay heat in SNF. Currently, the long term widely accepted solution is the geological disposal. At the same time, advanced technologies like Partitioning and Transmutation (P&T) offer promising solutions to reduce SNF long-term radiotoxicity. While most transmutation strategies rely on neutron fluxes, in this study the adoption of photon beam to induce photonuclear reactions in SNF is investigated, without depending on neutron based systems. In particular, the study focuses on the probability of inducing transmutations and fissions on MAs, by leveraging the Giant Dipole Resonance (GDR) region of photonuclear interactions. In the presented case study, the aim was to investigate the effect of a photon driven transmutation of minor actinides present in a spent fuel from SMR technology. The focus was mainly put on studying the physics of the system, analyzing the feasibility of reducing the inventory and radiotoxicity of the system by this method, without considering technological aspects and limitations.
@@
In 2020, the International Atomic Energy Agency (IAEA) reported that 172 Nuclear Power Plants (NPPs) have been permanently shut down and 20 have been fully decommissioned [1]. During the decommissioning and disposal phases, the main concern must be on the management of Spent Nuclear Fuel (SNF) due to its high and long-term radiotoxicity.
The composition of SNF varies depending on the reactor design and the fuel burning. Thermal reactors, which operate using a slower neutron spectrum, tend to generate more minor actinides (such as Americium, Curium, and Neptunium) than reactors based on a fast neutron spectrum. Indeed, at thermal neutron energies, the likelihood of neutron absorption is higher, leading to the production of transuranic elements. As a result, the composition of SNF from these reactors typically consists of approximately 94.5% uranium, 4.3% fission products, 1% plutonium (Pu), and 0.2% minor actinides (MAs) [2]111Ratios between elements in SNF depends on the burnup level at the discharge of the reactor. Reported values are representative of SNF at 50 MWd/kg.. Concerning the fast reactor fleet, the evaluation of typical spent fuel data is not trivial, since most of the fast technology adopts a closed fuel cycle with processed fuel, mixing MAs with fresh fuel [3].
@@
Regarding transmutation strategies, the vast majority of designs discussed in the literature focus on neutron-based transmutation, which rely on high neutron fluxes along with a fast neutron spectrum. This creates a scenario where the main driver for the consumption of minor actinides (i.e., neutrons) also contributes to their production. Consequently, there are no viable alternatives for SNF generated by thermal reactors. This limitation has prompted researchers to investigate non-neutron transmutation strategies, including facilities that generate high-energy particles other than neutrons. Recent advancements in photonics and ion accelerators present promising opportunities, enabling the production of energetic non-neutron particles at national lab user facilities or university accelerators. In addition, the study of alternative systems to fast
reactors for transmutation is attractive because of the lower engineering complexity that an accelerator requires compared to a nuclear plant. However, it is fair to point out that the difficulty in building a nuclear plant is repaid by the energy production, whereas in an accelerator-based scenario, there is no energy production but only consumption.
@@
Bremsstrahlung is the simplest and most widely used method for generating high photon fluxes. This process, induced by the deceleration of intense electron beams in high-Z materials like tungsten, lead, tantalum, or depleted uranium, generates photons with a continuous energy distribution. Materials with higher proton numbers are particularly effective as converters due to their lower critical energy thresholds. Bremsstrahlung facilities are invaluable for generating intense photon beams with straightforward technology, finding applications in experiments such as the search for fission isomers and the production of neutron rich radioactive ion beams [20]. Despite these advantages, Bremsstrahlung sources have significant limitations, including broad energy spectra, large background signals, and lack of polarization. However, these drawbacks are negligible in programs where the focus is on beam intensity rather than energy precision [21].
In contrast, recent advancements have enabled the production of high-intensity, quasi-monoenergetic gamma beams through laser Compton backscattering (LCB). This process, illustrated in Fig. 1, involves the interaction of laser photons (energy E=h) with GeV-level accelerated electrons: the photons gain energy from the relativistic recoil interaction, with the resulting energy scaling quadratically with the Lorentz factor of the electrons. In this manner, gamma-ray beams in the GDR energy region (10–20 MeV) can be achieved starting with laser photons of relatively low energy. The ability to produce photon beams with a narrow energy bandwidth, strong polarization, and a high degree of tuning makes this method particularly attractive for experiments requiring precise energy control, e.g. for precise evaluation of the GDR region cross section of target nuclei.
@@
Based on the above considerations, in this work a methodology to investigate the photon-driven transmutation of key MAs is developed, having focus on the fundamental physical processes involved in the MAs transmutation and proposing a new strategy for radiotoxicity reduction without relying on neutrons as the primary driver. Specifically, we explore photonuclear interactions for inducing transmutation and even fission by leveraging the Giant Dipole Resonance (GDR) in nuclei. The focus is on photo-absorption reactions, such as photocapture and photofission, triggered by high-energy photons, that have the maximum probability of occurring within the GDR region. To have a complete indication of the system’s behaviour concerning the source, both photons emitted by Bremsstrahlung and LCB are considered. The considered MAs are a subset of the radioactive waste produced by a fuel assembly from the NuScale reactor design [22], one of the most advanced and commercially ready SMR technologies. Specifically, the content of actinides present at the reactor discharge is considered separable into two groups: elements that can be recycled for power generation purposes and those that cannot. The former, such as U and Pu, for example, are not considered to be subject to irradiation, since they can be used in so-called MOX fuels; the latter, such as Np, Am and Cm, which are the main contributors to radiotoxicity, are subject to irradiation. Due to the transmutations that can occur on Np, Am and Cm, this analysis considers the build-up of up to 53 nuclides. It is important to emphasize that the purpose of this work is not to present a general and industrially competitive strategy, but rather to test the feasibility on the purely physical side of minor actinide transmutations from photons, also identifying the order of magnitude of the fluxes needed to reduce the system radiotoxicity. For this reason, similar to the work of Willat et al. [23] the discussion regarding the study of the costs of producing these fluxes is deliberately not considered in this paper: the initial average fluxes in the irradiated volume were taken equal to those reported by Matsumoto et al. [18], who performed a similar work, but focusing on some LLFPs.
@@ 2.1 Cross section data
Photo-nuclear reactions are fundamental processes in which atomic nuclei interact with high-energy photons, typically high-energy X-rays, resulting in various nuclear transformations. These reactions include photo-disintegration, where a nucleus absorbs a photon and subsequently emits nucleons such as neutrons or protons, and photo-excitation, where the nucleus is excited to a higher energy state [21]. These latter two phenomena appear to have different probability of occurrence depending on the energy of the impinging photon on the nucleus. This probability, encoded in the the photonuclear absorption cross section, can be divided in three main regions[24], depending on the photon energy. The GDR region ( 50 MeV), the Quasideuteron region ( 200 MeV) and the Photomeson region ( 200 MeV). The GDR region is interesting for requiring photon energies lower than 50 MeV and exhibiting high cross sections due to resonant effects. Also, for MAs, this region begins at approximately 5 MeV. More details about the photofission mechanism can be found in Balabanski et al.[21].
@@
| - | All 53 nuclides | - | |
| - | 232,237,239-240U, 232-236,238-244Np, 232-237,242-244Pu, 234-240,242-245Am, 240-245Cm | 238-241Pu, 233-236,238U, 237Np, 241Am | |
| - | 232-233,240-244Np, 232-235Pu, 234-239,245Am | 232-240U, 234-239Np, 236-244Pu, 240-244Am, 240-245Cm | |
| - | 232-233,240-244Np, 232-235Pu, 234-239,245Am | 232-240U, 234-239Np, 236-244Pu, 240-244Am, 240-245Cm | |
| - | 232-233,240-244Np, 232-235Pu, 234-239,245Am | 232-240U, 234-239Np, 236-244Pu, 240-244Am, 240-245Cm |
Finally, this work relies on a diverse array of nuclear data sources, given the lack of a single comprehensive database containing all relevant reaction information. The considered libraries are JENDL-5 [25], TENDL-2021 [26] and ENDF/B-VIII.0 [27]. Table 1 classifies the adopted nuclear data libraries for each considered reactions for the 53 isotopes. It is evident that the data were preferably taken from ENDF/B-VIII.0 and, when necessary, were completed with data taken from the other two libraries. A complete analysis should also consider the uncertainties related to each cross section and propagate the error in successive calculations. However, at the time this analisys was performed, the only nuclear library that contains covariance matrices is TENDL-2023. We plan to perform verification analysis relying on that nuclear library only, to also get the uncertainty propagation on the results.
@@ 2.2 Bateman equation solution
The information of the cross section alone is not sufficient to evaluate the transmutation capability of photons against neutrons. To do so, it is also crucial to compare the absolute values of reactions such as and for neutrons in the energy range above 0.5 MeV (see Figure 2). In this energy range, when neutron and photon fluxes are of equal intensity, the probability of neutron interactions is approximately three times higher than that of photon interactions. However, the actual intensities of these two fluxes differ significantly. The neutron flux is generated as secondary particles during photonuclear interactions and fission, while the photon flux originates from an accelerator, which produces a known and controllable beam. A more complete discussion about fluxes will be faced in Section 3, showing how to extract volume averaged energy fluxes (i.e., ) from the system and how the normalization was set in this work.
@@
Where represents the atomic density of nuclide in unit of atom/cm3.
Equation (1) appears to be linear and easily solved through the exponential matrix inversion only if the energy fluxes and are not time dependent. This hypothesis holds only if the change in atomic density is small relative to the initial condition; namely, a large change in atomic density induces a perturbation volume averaged fluxes, due to differences in attenuation.
For this reason, this work adopts a forward Euler scheme to take into account the modifications in the particle fluxes during time. This method is simple but effective: first, the time frame in which the equation will be solved is divided in substeps. During the first substep, the system of equations is solved with the spent fuel initial condition and with the initials volume averaged fluxes kept constant. At the end of the substep, the final composition is recorded and it is adopted for the re-evaluation of the fluxes. Then, the solution can proceed with the second substep, up to the end of the time frame. The presented algorithm is reported for clarity in Algorithm 1.
@@ 2.3 Radiotoxicity evaluation
Contents of radioactive isotopes in spent fuel can be assessed in terms of mass, activity, energy release or radiotoxicity. The latter figure of merit gives quantitative information on the radiobiological influence on the human body. Hazard from minor actinides is mainly due to alpha decay, which having long half-time, require strategies for their disposal or treatment.
@@
Also, we focused to the radioactivity content coming from MA only, without considering production/transmutation of LLFP that may build up during irradiation.
@@
where is the irradiated volume where actinides are uniformly distributed, in cm3, are the dose coefficients, tabulated within the ICRP 2012 report [28] and is the outcome concentration of the Bateman equations in at/cm3.
The system radiotoxicity is compared with the free evolution radiotoxicity , namely the Sievert emitted if the same mass of minor actinides used as the initial condition for this calculation (i.e., ) were to evolve without any particle flux acting as a driver. This concentration evolution has been evaluated with an independent Serpent simulation adopting the decay step mode, where Bateman equations are internally solved from the code, therefore obtaining the concentrations :
It is worth noticing that decay constants adopted for the evaluation of were taken from ENDF/B-VIII.0[29]. Both the quantities and have been normalised with respect to the natural uranium radiotoxicity , calculated as
@@ 3 Case study
The spent fuel inventory is primarily influenced by both the reactor design and the nominal power. The reactor design affects the composition of fresh fuel, including the types of enrichment and fissile materials present, as well as the evolution of the neutron energy spectrum. For instance, a reactor operating with a thermal spectrum tends to burn odd nuclei more quickly, while a reactor with a fast spectrum can effectively burn both odd and even nuclei. The nominal power, on the other hand, determines the rate at which the fuel is consumed, providing insight into the number of minor actinides generated through neutron captures over the reactor lifetime. The figure of merit that encapsulates fuel history is burnup (in MWd/kgU), which is proportional to the total energy produced by the reactor. A complete report with spent fuel inventory from different reactor technology and disposal choices could be found at [30].
The methodology presented in this paper is general, and its application is not restricted to the presented case only. However, we believe that it could be interesting for a general reader to learn the details of one of the most promising designs for the near future of energy generation from nuclear sources.
@@ 3.1 Initial condition: spent fuel composition
The initial conditions is based on the minor actinides present in the whole NuScale fuel, shown in Table 2. It is important to note that uranium and plutonium were deliberately excluded from this calculation, as their reuse in MOX fuels is widely expected in the literature[33]. The actinides considered in this work suffers from being not viable as nuclear fuel at present due to their poor neutronic performances, moreover they are responsible for a significant radiotoxicity level (i.e., 241Am and 237Np).
@@ 3.2 Flux evaluation in different scenarios
The interaction between a particle beam and the spent fuel is not trivial due to the high number of reaction channels available. Since this work uses a lumped approach, each isotope transmutation is driven by an average reaction rate, depending both by the microscopic cross section as well as by the average flux value. To have a precise estimation of the volume-averaged particle flux in a spent fuel sample irradiated by a particle beam, a set of numerical simulations are performed using the FLUKA code (version 2024.1.0) [34]. FLUKA is a Monte Carlo radiation transport code that offers high-precision modeling of particle interactions and transport in complex 3D geometries. It supports a wide range of particles (neutrons, photons, electrons, hadrons, heavy ions) across a broad energy spectrum, using continuous-energy cross sections and detailed physics models. FLUKA naturally handles coupled transport, energy deposition, and secondary particle production, making it ideal for radiation shielding, dosimetry, and activation analyses. Its flexible source definition, advanced variance reduction techniques, and extensive built-in scoring capabilities allow accurate and efficient solutions to complex radiation transfer problems in both stationary and time-dependent scenarios.
The irradiated sample is a cylinder having radius equal to 1 cm and height equal to 4 cm. This is chosen considering the attenuation of the primary beam and the secondary particles produced inside the spent fuel: the 12 cm3 cylinder delimits the zone where both the primary beam intensity and the secondary particle flux are reduced by a factor 1000. The volume is filled with a fictitious ceramic material MO2, where M stands for the isotopic mix shown in Tab. 2. The density has been set equal to 10.2 g/cm3, in the same order of magnitude of common oxides adopted in the nuclear field.
@@
The initial normalization of such fluxes is a crucial point in the presented analysis. The fluxes calculated by FLUKA are known despite a normalization factor F333For example, the FLUKA simulation evaluates the photon flux in the unit of , which depends by the source intensity of the accelerator. We considered different photon sources (i.e., bremsstrahlung and LCB) and our purpose is to compare the potentials of these sources and their effect on radiotoxicity by considering them at the same averaged flux volume. For this reason, the normalization factor was chosen by constraining the number of within the spent fuel sample. Namely, for each simulation:
Where is the volume-averaged flux and represents the beginning of the irradiation period. The value of Fs can also be adopted to normalize the neutron flux . This work considers four different flux scenarios values, i.e. 1E+17, 1E+18, 1E+19, 1E+20 . These values of flux are taken for reference from a similar work made by Matsumoto et al.[18]. It is important to highlight that the considered values were considered for having a comparison with the work of Matsumoto et al. and may not reflect realistic source intensities, which is not the purpose of this work. In particular, different ranges of fluxes are present nowadays for LCB and Bremmstrahlung facilities, the first ranging from 1E5 up to 1E11 , while the latter having maximum to 2E18 [21].
Figure 6 reports the scored photon and neutron fluxes with the initial spent fuel composition in the different simulations: with the adoption of the USRTRACK card, the number of tracks of photons and neutrons are recorded in an equi-spaced energy mesh. Then, the shapes are normalized according to Eqn. (5). It is interesting to note few differences in fluxes arising from Bremsstrahlung for diverse Tantalum thicknesses (cases A) and for no Tantalum converter (case B), whereas an LCB source (case C) is able to reproduce quasi-monoenergetic photons in the target volume and a more intense neutron field within the spent fuel.